OpenStudy (anonymous):

if v(t)= 50 + ce^0.196t by putting different values of c what is the process of getting the values for v(t)?

OpenStudy (anonymous):

we can also put random values for t

OpenStudy (anonymous):

all I want to know that how to solve e

OpenStudy (anonymous):

do we have to just puglin it's value 2.71828 ?

OpenStudy (turingtest):

I'm not sure I understand the question... for e you just use the number 2.71828... the value of c you have will change the shape of your graph, and each corresponds to a different integral curve

OpenStudy (turingtest):

all the different blue lines here (above ex 1) are the function v with different values of C so the exact function depends on the value of C, which is found by applying the initial conditions (if there are any)

OpenStudy (anonymous):

at http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx right after the example number 1 the author has sketched a graph, he claims that he made this graph by putting different values of c in v(t)=50+ce^-0.196t , for example if I want to put c=20 then how will I get the value?

OpenStudy (turingtest):

forgot to post the link http://tutorial.math.lamar.edu/Classes/DE/DirectionFields.aspx

OpenStudy (turingtest):

what value? you get a function, not a single value

OpenStudy (turingtest):

put C=20 and you get v(t)=50+20e^-0.196t which is just a function of t

OpenStudy (anonymous):

yeah the value of the function "v(t)"

OpenStudy (turingtest):

at what time?

OpenStudy (anonymous):

v(t)=50+20e^-0.196t how to solve it? :|

OpenStudy (anonymous):

lets say 1

OpenStudy (turingtest):

it's solved, just plug in whatever value for t you need if t is in minutes and you want the velocity at 5 min, then plug in t=5 and you get your answer

OpenStudy (turingtest):

let's say 1?

OpenStudy (anonymous):

I mean numerically or on calculator what value of e should I plug

OpenStudy (anonymous):

2.71828

OpenStudy (turingtest):

if we let C=20 and t=1 we have\[v(1)=50+20e^{-0.196}\]yes, if your calculator does not have e just put in however many digits you want for the number 2.718281828459045...

OpenStudy (anonymous):

at c = 20 & t=1 I get 50.4456365691809 whereas at c20 i should get 70 isn't?

OpenStudy (anonymous):

ooopps I am forgetting to add 50

OpenStudy (anonymous):

got it now

OpenStudy (anonymous):

thanks @TuringTest

OpenStudy (turingtest):

very welcome, as usual :D